XYZ Company produces three products, A, B, and C. XYZ's plant capacity is limited to 200,000 machine hours per year. The following information is available for planning purposes:

Product A Product B Product C

Demand for Next Year

100,000 units 200,000 150,000
Selling Price Per Unit

18

28 23
Variable costs per unit 15 20 21

It is known that it takes 0.50 machine hours to produce one unit of Product A; 0.80 machine hours to produce one unit of Product B; and 0.40 machine hours to produce one unit of Product C. Calculate the number of units of Product C that XYZ Company should produce in order to maximize its net income.
Answer #1

Solution:

13 14 15 16 Computation of contribution margin per Machine hour 17 А 18 Selling price per unit $18.00 $28.00 19 less: variabl

According to the ranking, first demand for product B and then A will be fully produced and then Product C will be produced if machine hours left.

Machine hours used to fulfil Product B 's demand = 200000*0.80 = 160,000 machine hours

Machine hour needed for product A = 100000*0.50 = 50,000 machine hours

Now, since machine hours exceeded the available machine hours, product C cannot be produced.

Hence, unit of product C = 0